OpenStudy (anonymous):

2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (anonymous):

To solve this, which method do you prefer? substitution or elimination?

OpenStudy (anonymous):

This system requires graphing, but I don't have any graph paper. I did get this though 3x-2y=11 > 2y = -3 +11 and 2x + 3y =3 >3y = -2x + 3

OpenStudy (anonymous):

Well, you did some parts right, but lets do it again: 3x-2y=11 -2y=-3x+11 y=3/2x-11/2 Did you notice where you made the mistake? and the other one: 2x+3y=2 3y=-2x+2 y=-2/3x+2/3 Now before I graph it, did you notice your mistake?

OpenStudy (anonymous):

I left out the x after -3?

OpenStudy (anonymous):

Yes, and in my second equation, I also made a mistake, it should be y=-2/3x+1

OpenStudy (anonymous):

okay~

OpenStudy (anonymous):

So I just graphed it, and it intersects at 3,-1. Which is your answer. Sometimes, you just can also put the options in the equations, and see which option makes it true.

OpenStudy (pixiedust1):

Are you still there?

OpenStudy (anonymous):

Me?

OpenStudy (pixiedust1):

No, the asker

OpenStudy (anonymous):

Yes?

OpenStudy (pixiedust1):

I have an easier way to solve this.

OpenStudy (pixiedust1):

MUCH easier. :D

OpenStudy (anonymous):

Substitution?

OpenStudy (pixiedust1):

Yuppers. Just substitute them all in.

OpenStudy (pixiedust1):

I'm sure you know how to do that

OpenStudy (anonymous):

Okay,I'll try that next time. (And yes, I do, I just wanted to make sure I know how to do it this way too)

OpenStudy (pixiedust1):

That's terrific! That's what a smart student does! Here is a badge for you. :D

OpenStudy (anonymous):

Oh, thank you~

OpenStudy (anonymous):

find the y first